Verify mean value theorem, if f(x) =x2-4x-3 in the interval [a,b] where a= 1 and b= 4 x


1
Expert's answer
2021-10-27T12:26:43-0400

"f(x)=x^2-4x-3" is continuous on "[1,4]" as polynomial.


"f(x)=x^2-4x-3" is differentiable on "(1,4)" as polynomial.


Therefore, by the Mean Value Theorem, there is a number "c" in "(1, 4)" such that


"f(4)-f(1)=f'(c)(4-3)"

"f(4)=4^2-4(4)-3=-3"

"f(1)=1^2-4(1)-3=-6"

"f'(x)=2x-4"

"f'(c)=2c-4"

Substitute


"-3-(-6)=(2c-4)(4-1)"

"2c-4=1"

"c=\dfrac{5}{2}"

"1<\dfrac{5}{2}<4=>c\in (1, 4)"

Hence the hypotheses of the Mean Value Theorem are satisfied.